OpenStudy (anonymous):

[8.04] Part 1: Jameel and Sarah are discussing how to factor 3b2 + 8b + 4. Jameel feels this trinomial is prime because he cannot find factors of 4 that have a sum of 8. Sarah says he is incorrect and that it is factorable. Using complete sentences, provide a convincing argument explaining who is correct and why. If this trinomial is factorable, factor it showing all work and explain your steps. Part 2: Create your own prime trinomial in the form ax2 + bx + c. Using complete sentences, explain how you know it is prime.

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

bro ?

OpenStudy (anonymous):

are you doing it ?

OpenStudy (anonymous):

Part 1: Since leading coefficient is not 1 (it's 3), then we need to find factors of 3 * 4 = 12 that have a sum of 8. There are such factors: 2 and 6. Now replace 8b with 6b + 2b, then factor two terms at a time: 3b^2 + 8b + 4 = (3b^2 + 6b) + (2b + 4) = 3b (b + 2) + 2 (b + 2) = (b + 2) (3b + 2) So Jameel is wrong - trinomial is not prime - it can be factored. Here's the factored trinomial: (3b + 2)(b + 2). You also have to find factors of the squared number too, which in this case is not 1 and 1 like in other trinomials, but rather something different. Prime numbers are the easiest to work with because there's only one combination for the factors. Sarah is right . Part 2: To create a prime trinomial, set a as 1 and then find a b and c that don't factor like b = 5, c = 2.